# Performance and Storage Analysis of CRYSTALS-Kyber (ML-KEM) as a Post-Quantum Replacement for RSA and ECC

### v1

@@ 5.2 Storage benchmarks
Conclusion
The benchmarking results demonstrate CRYSTALS-Kyber’s decisive computational advantages over classical cryptographic schemes across both x86_64 and ARM64 architectures. These gains reflect Kyber’s natural alignment with modern CPU features; its implementation leverages vectorization instructions such as AVX2 to accelerate core lattice-based computations. Kyber operates 2.7–3$\times$ faster than ECC and 3,400$\times$–20,500$\times$ faster than RSA in key generation due to its efficient lattice-based polynomial arithmetic. For shared secret derivation, Kyber maintains 41–72$\times$ speed advantages over ECC and 1,600$\times$–3,200$\times$ advantages over RSA, effectively eliminating RSA’s critical bottleneck in receiver-side operations. These performance gains are consistent across architectures, with ARM64 partially mitigating the overheads of classical schemes without altering the fundamental performance hierarchy. The sole trade-off emerges in storage requirements, where Kyber’s 1,121-byte hybrid payload exceeds RSA’s 993 bytes and ECC’s minimal 30-byte overhead, reflecting the inherent tension between quantum resistance and bandwidth efficiency. Collectively, these results establish Kyber as delivering NIST-standardized quantum resistance without computational penalties.

### v2

@@ 5.2 Storage benchmarks
The added overhead refers to the data that an algorithm adds to the existing data (plaintext length) to encrypt it, and in some cases, to include the key transport, for example, with Kyber, is: plaintext message length + Kyber’s overhead + Rust’s serialization bytes. The storage overhead analysis reveals a critical trade-off between quantum resistance and bandwidth efficiency. While the Poly1305 authentication nonce contributes a constant 30-byte overhead across all schemes, the key transport mechanism diverges significantly: SECP384R1 (ECC) requires zero key transport overhead due to its elliptic curve key agreement protocol, resulting in a minimal total added length of 30 bytes. In contrast, Kyber and RSA introduce substantial key transport payloads (1,091 bytes and 963 bytes, respectively), resulting in total overheads of 1,121 bytes for Kyber and 993 bytes for RSA. This establishes ECC as the most bandwidth-efficient scheme—advantageous in constrained environments, such as IoT—but its quantum vulnerability remains an existential risk. Kyber’s $13.5\%$ higher transport overhead than RSA is justified by its quantum resistance and superior computational efficiency (see Section 5.1), positioning it as the optimal choice for future-proof systems where bandwidth permits.
Conclusion